how do you solve: for her workout, Sarah walks north at the rate of 3mph and returns at the rate of 4mph. how many miles did she walk if the round trip took 3.5hours?
Answer by stanbon(75887)

Sarah walks north at the rate of 3mph and returns at the rate of 4mph. how many miles did she walk if the round trip took 3.5hours?
------------------------
North DATA:
rate = 3 mph ; time = x hrs ; distance = rt = 3x miles
---------------------------------
Return DATA:
rate = 4 mph ; time = 3.5-x hrs ; distance = rt = 4(3.5-x) miles
----------------------------------------------------------------------
Equation:
distance = distance
3x = 4(3.5-x)
3x = 14-4x
7x = 14
x = 2 hrs (North time)
---
North distance = 3x = 3*2 = 6 miles
Return distance = 4(3.5-2) = 6 miles
----
Round-trip = 6 miles
===========================
Cheers,
Stan H.